OpenStudy (anonymous):

Find the smallest angle with sides 2m, 3m, 4m.

OpenStudy (anonymous):

The angle opposite to the smallest side is smallest angle in a trianglr

OpenStudy (anonymous):

I mean triangle

OpenStudy (anonymous):

saura is right.

OpenStudy (anonymous):

and we will use cosine for this right?

OpenStudy (anonymous):

A greater angle of a triangle is opposite a greater side.

OpenStudy (anonymous):

2^2 = 4^2 + 3^2 2(4)(3) (cos x)

OpenStudy (anonymous):

I see can someone show me the cosine step, I keep messing up the BEDMAS

OpenStudy (anonymous):

Well if you have to find the exact angle. Haven't gotten there yet.

OpenStudy (anonymous):

YA use cosine rule if u want to find the exact value of the angle

OpenStudy (anonymous):

well can you show that for me, I tried a lot but failed each time. My steps just doesn't help me get the answer.

OpenStudy (anonymous):

U want to find the value of smallest angle?

OpenStudy (anonymous):

yes. that's what the question asks.

OpenStudy (anonymous):

let x be the smallest angle

OpenStudy (anonymous):

then, (2m)^2=(3m)^2+(4m)^2-2*3m*4m*cosx

OpenStudy (anonymous):

Now, solve for x

OpenStudy (anonymous):

wait why is there an m there? Regardless, 4m^2 = 9m^2 + 16m^2 - 24m*cosx

OpenStudy (anonymous):

did i do that right?

OpenStudy (anonymous):

(2m)^2=(3m)^2+(4m)^2-2*3m*4m*cosx 4m^2=9m^2+16m^2-24m^2 cosx

OpenStudy (anonymous):

ohhhhh 4m^2 -25m^2 = -24m^2 cosx -21m^2/-24 = cosx 7/8 = cos x inverse cos(7/8) = x 28.95 = x

OpenStudy (anonymous):

thankss :)

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

:)yw